OpenStudy (anonymous):

Which of the following is a solution of x2 + 2x + 4?

OpenStudy (anonymous):

@amistre64

OpenStudy (gorv):

x^2+2*x+4 factorise it

OpenStudy (anonymous):

(x+2)(x-2)

OpenStudy (anonymous):

x=-2 and x=2

OpenStudy (anonymous):

but my answers can only be A-\[-2+2i \sqrt{3}\] B-\[2+2i \sqrt{3}\] C-\[1-1i \sqrt{3}\] D-\[-1+1i \sqrt{3}\]

OpenStudy (anonymous):

@E.ali Okay I get what you're saying so far I think

OpenStudy (gorv):

sorry factor will be x=-1(+-)i*sqrt3

OpenStudy (anonymous):

Thank you!